OpenStudy (anonymous):

White Silver Black Grey Total Car: 8 9 3 4 24 Truck: 2 4 7 4 17 SUV: 3 5 8 3 19 Total: 13 18 18 11 60 Let event A be defined as a randomly selected vehicle being silver or black. Let event B be defined as a randomly selected vehicle being a car or a truck. Find P(~ B|A) A.1/10 B.13/60 C.1/4 D.13/36

OpenStudy (nikato):

So how many cars aren't black or silver

OpenStudy (anonymous):

24

OpenStudy (nikato):

And how many cars and trucks?

OpenStudy (anonymous):

41

OpenStudy (nikato):

Nvm. I think I did something wrong or misread

OpenStudy (anonymous):

is that P(~B | A) or P(~ (B|A) ) ?

OpenStudy (anonymous):

P(~ (B|A)

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

i would just guess A , Yolo.

OpenStudy (anonymous):

P(~ (B|A) ) = 1 - P(B|A)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

1 - 23/36 = 13/36 so D